X company is not a highly leveraged company with 400 million shares, trading at $75/share and $3 billion in debt (in market value terms) outstanding. The firm has a A3 rating with a default spread of 1.1% over the riskfree rate.

Estimate the pre-tax cost of debt.
a 1.10%
b 1.65%
c 2.75%
d 3.85%
e 1.82%
Answer #1

Pre tax cost of debt = Risk free rate + Spread

= 2.38+1.1

= 3.48

Closest Answer being 3.85 we can assume risk free return to be 2.75

hence Cost of debt = 2.75+1.1 = 3.85

Note 1

The risk-free rate of return is the theoretical rate of return of an investment with zero risk.

In theory, the risk-free rate is the minimum return an investor expects for any investment because he will not accept additional risk unless the potential rate of return is greater than the risk-free rate.

In practice, however, the risk-free rate does not exist because even the safest investments carry a very small amount of risk. Thus, the interest rate on a three-month U.S. Treasury bill is often used as the risk-free rate for U.S.-based investors.

3 Month Treasury Bill Rate is at 2.38% as on 8 may 2019